OpenStudy (anonymous):

A company offers you a job with two choices of salarys: A: S= 40,000 (1.025)^t B: S=1500t+40,000. How many years until option A overtakes option B?

OpenStudy (anonymous):

T= years, S= salary

OpenStudy (ganpat):

A: S= 40,000 (1.025)^t B: S=1500t+40,000. equate them.. 40,000 (1.025)^t = 1500t+40,000 40,000 (1.025)^t = 1500t+40,000 i cn take u till this far :-p..

OpenStudy (ganpat):

@ganeshie8 , @mathslover , @ParthKohli .. guys, need help..

OpenStudy (anonymous):

Yeah I get to there too

OpenStudy (anonymous):

Don't know how to solve

OpenStudy (ganpat):

neither do i :-p.. sorry !!

OpenStudy (anonymous):

Thanks anyway